jillsacc:

The table and the graph below each show a different relationship between the same two variables, x and y: A table with two columns and 5 rows is shown. The column head for the left column is x, and the column head for the right column is y. The row entries in the table are 4 comma 120 and 5 comma 150 and 6 comma 180 and 7 comma 210. On the right of this table is a graph. The x-axis values are from 0 to 10 in increments of 2 for each grid line. The y-axis values on the graph are from 0 to 550 in increments of 110 for each grid line. A line passing through the ordered pairs 2 comma 110 and 4 comma 220 and 6 comma 330 and 8 comma 440 is drawn. How much more would the value of y be on the graph than its value in the table when x = 12? 150 300 450 600

SmokeyBrown:

Hi and welcome to QuestionCove! The table describes a line with the equation y1 = 30x1 The graph depicts a line with the equation y2 = 55x2 We can compare the value of y for each of these lines when x=12: y1 = 30*12 = 360 y2 = 55*12 = 660 So, how much more is the value of y on the graph compared to y on the table when x is 12? Thank you for your very detailed description of the question; if you find it easier, you may also consider attaching a screenshot as a comment on your post in order to share information in the future :)